OpenStudy (anonymous):

Part 1: Solve the inequality -3x + 4 <= x + 8 and show all of your work. Part 2: Use complete sentences to describe the graph of the solution from Part 1.

OpenStudy (anonymous):

-3x + 4 <= x + 8 4 <= 3x + 8 -4 <= 3x -4/3 = x

OpenStudy (anonymous):